Find the slope from the following pairs of points:
(5,-2)&(9,2) (1,-5)&(3,1) (0,6)&(2,6)

Answer by checkley71(8403)

(2+2)/(9-5) OR 4/4=1
(1+5)/(3-1) OR 6/2=3
(6-6)/(2-0) OR 0/2 OR STRIGHT VERTICLE LINE THROUGH Y=6